# Lectures on Gauge theories and Many-Body systems

### v1

@@ 3.2 $2d$ Yang–Mills theory
$\sqrt{-1}A_{t}\mapsto g^{-1}\sqrt{-1}A_{t}g-g^{-1}\dot{g},$
@@
$S=\int\mathrm{Tr}\big(P[\partial_{t}-\sqrt{-1}A_{t},X]-\frac{1}{2}P^{2}+(\partial_{t}-\sqrt{-1}A_{t})J\big)dt,$
@@
$J=\sqrt{-1}(z\otimes z^{\dagger}-\nu 1_{N}).$
Gauge theory motivates replacing $X$ to $\sqrt{-1}\partial_{x}+A_{x}$. Such a generalization turns the connection $d-\sqrt{-1}A_{t}dt$ on $\mathbb{R}\ni t$ into the connection $D=d-\sqrt{-1}A_{t}dt-\sqrt{-1}A_{x}dx$ on $\mathbb{R}\times\mathbb{R}\ni(t,x)$ (or $\mathbb{R}\times S^{1}$). Analogous object in electrodynamics is called vector potential. In this way $\partial_{t}-\sqrt{-1}A_{t},X]$ becomes a curvature form $\sqrt{-1}F_{tx}^{A}=\sqrt{-1}[\partial_{t}-\sqrt{-1}A_{t},\partial_{x}-\sqrt{-1}A_{x}]$. Being a conjugated variable to the $X$ the quantity $P$ must be replaced with a function $E(x)$. Its electrodynamical analogue is electric field. We want the term $J$ to impose some condition on $A_{x}$ at single point ($x=0$) only (we will see later that this is reasonable), so let us replace it by $J\delta(x)dx$. The new action thus has the form
$S_{YM}=\int\mathrm{Tr}(\sqrt{-1}EF^{A})-\frac{1}{2}\mathrm{Tr}(E^{2})dxdt+\mathrm{Tr}(D_{t}^{A}J)\delta(x)dxdt.$
This is the $2d$ Yang–Mills theory ([Migdal:1975]) in the first-order formalism. Let us rewrite it in the following way
$S_{YM}=\int\mathrm{Tr}\big(-\frac{1}{2}E^{2}+D_{t}(\sqrt{-1}D_{x}E-J\delta(x))\big)dxdt.$
@@ 3.3 Hamiltonian description
and impose the constraint $\mu_{\infty}\overset{def}{=}\sqrt{-1}D_{x}E-J\delta(x)=0$ (Gauss law).
@@
A new phase space formed by $(E(x),\sqrt{-1}\partial_{x}+A_{x}(x),z)$ is
@@ 3.4 Calogero–Moser–Sutherland system
Analogously to the previous case we need to find a nice representative in every orbit of $\mathcal{G}$. Every $u(N)$-connection $A_{x}$ on $S^{1}$ defines a parallel transport along the circle the monodromy matrix $M\in U(N)$. Gauge invariant eigenvalues $e^{2\pi\sqrt{-1}x_{1}},...,e^{2\pi\sqrt{-1}x_{N}}$ of $M$ provide natural set $(x_{1},...,x_{N})$ of functions on $\mathcal{P}_{\infty}$ well defined on $\mathcal{P}_{\infty}/\mkern-5.0mu/\mathcal{G}$ up to permutations and shifts:
$g=\mathrm{diag}(e^{\sqrt{-1}n_{i}x}):(x_{1},...,x_{N})\mapsto(x_{1}+n_{1},...,x_{N}+n_{N}),\penalty 10000\ n_{i}\in\mathbb{Z}.$
@@
Note that such transformations form an affine Weyl group $S_{N}\ltimes\mathbb{Z}^{N}$ of $\widehat{u(N)}$. This is the analogue of the Weyl group $S_{N}$ of $u(N)$ acting by permutations on eigenvalues of $X$ in the finite dimensional case.
Using the Floquet-Lyapunov theory ([Fl], [Lyap])one can prove that every $A_{x}$ can be gauged to the constant matrix of the form $A_{x}=\mathrm{diag}(x_{1},...,x_{N})$. The remaining symmetry is
$\mathrm{Stab}_{\mathcal{G}}(\mathrm{diag}(x_{1},...,x_{N}))=\{g(x)=\mathrm{diag}(e^{\sqrt{-1}\theta_{1}},...,e^{\sqrt{-1}\theta_{N}}):\theta_{i}=\mathrm{const}\in\mathbb{R})\}=U(1)^{\times N}.$
@@
Note that if $z_{i}$ were functions of $x$, the reduced phase space $\mathcal{P}_{S}$ would be infinite dimensional (since the remaining gauge symmetry $U(1)^{\times N}$ has finite dimension). That is why in our generalization we left $z_{i}$ as constants rather then making them fields.
@@
$E_{ij}(x)=\begin{dcases}p_{i},\;\;i=j\\ \frac{\sqrt{-1}\nu e^{-\pi\sqrt{-1}(x_{i}-x_{j})}}{2\sin(\pi(x_{i}-x_{j}))}e^{\sqrt{-1}(x_{i}-x_{j})x},\;\;i\neq j\end{dcases}$
@@
This is the Hamiltonian of Calogero–Moser–Sutherland system (also called the trigonometric CM system) ([Sutherland1971], [Moser1975]).
The construction above is an example of the promised direct correspondence between classical integrable systems and gauge theories. Here is a list of another examples:
@@
    Relativistic rational Ruijsenaars–Schneider ([Ruijsenaars:1986vq], [Ruijsenaars:1986pp]), [Ruijsenaars:1988pv], [RuijSim1990], [Ruijsenaars:1994gc])
@@
    Trigonometric Ruijsenaars–Schneider corresponds to $3d$ Chern–Simons ([Witten:1988hc])
@@
    Elliptic Ruijsenaars–Schneider corresponds to hybrid topological-holomorphic $4d$ Chern–Simons ([Nekrasov1996],[costello2013])
In the following the Elliptic Calogero–Moser will be considered.
@@ 4 Elliptic Calogero–Moser system
Elliptic Calogero–Moser system ([Calogero:1975]) is a Hamiltonian system containing a triple
@@
One can also consider real valued $x_{i}$ and $p_{i}$. Analogously to the usual CM this system describes motion of particles interacting with a slightly more sophisticated potential. However allowing positions and momenta to be complex provides an interesting feature: points can swip without gaining infinite energy. This circumstance prompts us to consider an extended phase space.
@@ 4.1 Compactification of the phase space
Instead of giving a definition we illustrate an idea for special case of $N=2$ points. For simplicity we restrict on a center-of-mass frame
@@ 4.2 Around fixed points
$\begin{cases}X&=\zeta x\\ Y&=\zeta(y+\sqrt{-1}z)\\ Z&=-\zeta(y-\sqrt{-1}z)\end{cases}$
@@
$\omega^{\mathbb{C}}=dz_{1}\wedge dz_{2}=\frac{dY\wedge dZ}{4X}=\sqrt{-1}\zeta\frac{dy\wedge dz}{2x}.$
@@
$\langle[\omega^{\mathbb{C}}],[D]\rangle=\langle[\omega^{\mathbb{C}}],[S^{2}]\rangle=\int\limits_{S^{2}}\omega^{\mathbb{C}}=\sqrt{-1}\zeta\int\limits_{S^{2}}\frac{dy\wedge dz}{2x}=\sqrt{-1}\zeta\int\limits_{S^{2}}\mathrm{Vol}_{S^{2}}=4\pi\sqrt{-1}\zeta.$
@@ 4.3 Phase space as an algebraic variety
5 Complexification of tCM – 2d Yang–Mills correspondence
In the following we will describe the complex analogue of infinite dimension Hamiltonian system previously considered. The symplectic reduction in this setting will reproduce the elliptic CM system. By complexification we mean passage from left to the right in the following table.
@@ 5 Complexification of tCM – 2d Yang–Mills correspondence
$(g^{-1}Eg,g^{-1}A_{x}g-\sqrt{-1}g^{-1}\partial_{x}g,g^{-1}(0)z)$ $(g^{-1}Eg,g^{-1}\bar{A}g-g^{-1}\partial_{\bar{z}}g,g^{-1}(0)v,g(0)w)$
@@
$J=\sqrt{-1}(z\otimes z^{\dagger}-\nu 1_{N})$ $J^{\mathbb{C}}=\sqrt{-1}(\nu 1_{N}-v\otimes w^{t})$
@@
([Atiyah1957VectorBO]) General rank $N$ complex vector bundle of degree 0 on $C_{\tau}$ is isomorphic to $\bigoplus\limits_{i=1}^{N}L_{i}$, where each $L_{i}$ is a degree 0 line bundle.
@@ 5.1 Complex gauging
$g=\mathrm{diag}(e^{\frac{2\pi\sqrt{-1}}{\tau-\bar{\tau}}((z-\bar{z})n_{i}+(z\bar{\tau}-\bar{z}\tau)m_{i})}):x_{i}\mapsto x_{i}+m_{i}+n_{i}\tau.$
@@ 6 Gauge theories and measures on partitions
Supersymmetric gauge theories in four dimensions, and five and six dimensional theories compactified to four dimensions, produce a class of complexified statistical models. The random variables of these models are collections of Young diagrams, also known as colored partitions [Nekrasov:2002qd, Nekrasov2015BPSCFTCN]. More specifically, this is a result of a two-step localization procedure. First of all, it is well-known since the interpretation of M. Atiyah and L. Jeffrey of E. Witten’s topological quantum field theory intepretation [Witten:1988ze] of Donaldson theory[Donaldson:1990kn] , path integral of topologically twisted ${\mathcal{N}}=2$ supersymmetric gauge theory, for specific correlation functions, can be interpreted as an integral of equivariant differential forms (albeit in an infinite-dimensional setting). This interpretation has been generalized to theories with matter hypermultiplets [Vafa:1994tf, Labastida:1996tz, Moore:1997pc, Losev:1997tp]. Whenever such interpretation holds, path integral localizes to a finite dimensional integral over the moduli space of instatons, i.e. solutions to the $F^{+}_{A}=0$ equations modulo suitable gauge group. This localization is an infinite-dimensional version of the equivalence of different representatives of Euler class of a vector bundle: Poincare dual to the zero locus of generic transverse section, and the curvature Gauss-Bonnet type form. The next step undertaken in [Nekrasov:2002qd] is to deform the theory in such a way that the path integral measure becomes the equivariant differential form with respect to both the local gauge symmetries and the global symmetries consisting of the rotations of ${\mathbb{R}}^{4}$ (more generally, isometries of the four-manifold on which one studies the theory), the rotations of gauge frame and possibly rotations of the spaces of multiplicities of matter fields. The question of framing arises when one is working on ${\mathbb{R}}^{4}$ or other non-compact manifolds. We need to impose the boundary conditions both on gauge fields and gauge transformations one divides by. As a result, the gauge group $\mathcal{G}$$_{\infty}$ is a proper normal subgroup in the group $\mathcal{G}$} of all gauge transformations defined on some compactification of ${\mathbb{R}}^{4}$, e.g. $S^{4}={\mathbb{R}}^{4}\cup{\infty}$. The quotient ${\mathcal{G}}/{\mathcal{G}}_{\infty}$ is the group of frame rotations. The equivariant parameters $\bf a$ for this group are identified with the vacuum expectation value of the adjoint-valued complex Higgs field, the scalar in the vector multiplet of ${\mathcal{N}}=2$ supersymmetry:
@@
The way path integral localizes onto the instanton configurations and then onto the set of special instanton configurations has been described in literature on many occasions, see [Pestun:2016zxk] for a review. Using the localization in equivariant cohomology we express the integral as a sum over the corresponding fixed points. The subtlety is that the fixed points do not belong to the actual moduli space of smooth supersymmetric gauge field configurations, such as instantons. Rather, they are parked somewhere on the compactified moduli space. The points of the compactified moduli space no longer parametrize the solutions to the instanton equations on the smooth four dimensional spacetime. They can be interpreted in algebraic geometry as parametrizing torsion free sheaves on ${\mathbb{C}\mathbb{P}}^{2}$ – a compactification of ${\mathbb{C}}^{2}$, a complex plane, whose underlying four dimensional real space is the original spacetime. Every instanton uniquely corresponds to some locally free, and thereofore, torsion free sheaf. However, there are more general torsion free sheaves. A limit of a sequence of locally free sheaves could be a torsion free sheaf which is not locally free. Let us explain this in a little more detail.
@@
The abovementioned torsion free sheaves obey a weaker condition: if $f\in{\mathcal{O}}_{U}$ is such, that $fs=0$ for non-zero $s\in{\Gamma}({\mathcal{E}}|_{U})$, then $f=0$. However, given the space of local sections of a torsion free sheaf, one cannot, in general, find a $(0,1)$-gauge field $A_{\bar{i}}dz^{\bar{i}}$, so that (6.4) would hold. The issue is that all $\Psi\in{\Gamma}({\mathcal{E}}|_{U})^{\oplus N}$ may degenerate at some point, so that $A_{\bar{i}}$ defined from (6.8) would be singular there. A way to address this issue, by keeping the moduli interpretation of the (partly) compactified moduli space of instantons, is to work on a noncommutative ${\mathbb{R}}^{4}$ [Nekrasov:1998ss].
@@
in other words, form a partition. Equivariant localization procedure assigns a weight to each fixed point of the form (6.9), and thus defines a measure on a set of multipartitions. In sections 6.2-6.3 we introduce some basic notions relevant for partitions and discuss their properties. Then in sections 6.4-6.8 we give precise formulas for the measures on multipartitions arising from gauge theories in various examples.
@@
These notions are reviewed in [Nekrasov:1998ss]. Both (6.13) and (6.14) are invariant under the transformations

### v2

@@ 3.2 $2d$ Yang–Mills theory
${\mathrm{i}}A_{t}\mapsto g^{-1}{\mathrm{i}}A_{t}g-g^{-1}\dot{g},$
@@
$S=\int\mathrm{Tr}\big(P[\partial_{t}-{\mathrm{i}}A_{t},X]-\frac{1}{2}P^{2}+(\partial_{t}-{\mathrm{i}}A_{t})J\big)dt,$
@@
$J={\mathrm{i}}(z\otimes z^{\dagger}-\nu 1_{N}).$
Gauge theory motivates replacing $X$ to ${\mathrm{i}}\partial_{x}+A_{x}$. Such a generalization turns the connection $d-{\mathrm{i}}A_{t}dt$ on $\mathbb{R}\ni t$ into the connection $D=d-{\mathrm{i}}A_{t}dt-{\mathrm{i}}A_{x}dx$ on $\mathbb{R}\times\mathbb{R}\ni(t,x)$ (or $\mathbb{R}\times S^{1}$). Analogous object in electrodynamics is called vector potential. In this way $\partial_{t}-{\mathrm{i}}A_{t},X]$ becomes a curvature form ${\mathrm{i}}F_{tx}^{A}={\mathrm{i}}[\partial_{t}-{\mathrm{i}}A_{t},\partial_{x}-{\mathrm{i}}A_{x}]$. Being a conjugated variable to the $X$ the quantity $P$ must be replaced with a function $E(x)$. Its electrodynamical analogue is electric field. We want the term $J$ to impose some condition on $A_{x}$ at single point ($x=0$) only (we will see later that this is reasonable), so let us replace it by $J\delta(x)dx$. The new action thus has the form
$S_{YM}=\int\mathrm{Tr}({\mathrm{i}}EF^{A})-\frac{1}{2}\mathrm{Tr}(E^{2})dxdt+\mathrm{Tr}(D_{t}^{A}J)\delta(x)dxdt.$
This is the $2d$ Yang–Mills theory ([Mig75]) in the first-order formalism. Let us rewrite it in the following way
$S_{YM}=\int\mathrm{Tr}\big(-\frac{1}{2}E^{2}+D_{t}({\mathrm{i}}D_{x}E-J\delta(x))\big)dxdt.$
@@ 3.3 Hamiltonian description
and impose the constraint $\mu_{\infty}\overset{def}{=}{\mathrm{i}}D_{x}E-J\delta(x)=0$ (Gauss law).
@@
A new phase space formed by $(E(x),{\mathrm{i}}\partial_{x}+A_{x}(x),z)$ is
@@ 3.4 Calogero–Moser–Sutherland system
Analogously to the previous case we need to find a nice representative in every orbit of $\mathcal{G}$. Every $u(N)$-connection $A_{x}$ on $S^{1}$ defines a parallel transport along the circle the monodromy matrix $M\in U(N)$. Gauge invariant eigenvalues $e^{2\pi{\mathrm{i}}x_{1}},...,e^{2\pi{\mathrm{i}}x_{N}}$ of $M$ provide natural set $(x_{1},...,x_{N})$ of functions on $\mathcal{P}_{\infty}$ well defined on $\mathcal{P}_{\infty}/\mkern-5.0mu/\mathcal{G}$ up to permutations and shifts:
$g=\mathrm{diag}(e^{{\mathrm{i}}n_{i}x}):(x_{1},...,x_{N})\mapsto(x_{1}+n_{1},...,x_{N}+n_{N}),\penalty 10000\ n_{i}\in\mathbb{Z}.$
@@
Note that such transformations form an affine Weyl group $S_{N}\ltimes\mathbb{Z}^{N}$ of $\widehat{u(N)}$. This is the analogue of the Weyl group $S_{N}$ of $u(N)$ acting by permutations on eigenvalues of $X$ in the finite-dimensional case.
Using the Floquet–Lyapunov theory ([Flo83], [Maw05]) one can prove that every $A_{x}$ can be gauged to the constant matrix of the form $A_{x}=\mathrm{diag}(x_{1},...,x_{N})$. The remaining symmetry is
$\mathrm{Stab}_{\mathcal{G}}(\mathrm{diag}(x_{1},...,x_{N}))=\{g(x)=\mathrm{diag}(e^{{\mathrm{i}}\theta_{1}},...,e^{{\mathrm{i}}\theta_{N}}):\theta_{i}=\mathrm{const}\in\mathbb{R})\}=U(1)^{\times N}.$
@@
Note that if $z_{i}$ were functions of $x$, the reduced phase space $\mathcal{P}_{S}$ would be infinite-dimensional (since the remaining gauge symmetry $U(1)^{\times N}$ has finite dimension). That is why in our generalization we left $z_{i}$ as constants rather than making them fields.
@@
$E_{ij}(x)=\begin{dcases}p_{i},\;\;i=j\\ \frac{{\mathrm{i}}\nu e^{-\pi{\mathrm{i}}(x_{i}-x_{j})}}{2\sin(\pi(x_{i}-x_{j}))}e^{{\mathrm{i}}(x_{i}-x_{j})x},\;\;i\neq j\end{dcases}$
@@
This is the Hamiltonian of Calogero–Moser–Sutherland system (also called the trigonometric CM system) ([Sut71], [Mos75]), derived from $2d$ Yang-Mills theory, following [GorNek1].
The construction above is an example of the promised direct correspondence between classical integrable systems and gauge theories. Here is a list of further examples:
@@
    Relativistic rational Ruijsenaars–Schneider ([RS86], [Rui87]), [Rui88], [Rui90], [Rui94]) corresponds to the topological $2d$ Yang-Mills, perturbed by a Wilson loop operator [FGNR].
@@
    Trigonometric Ruijsenaars–Schneider corresponds to $3d$ Chern–Simons ([W88a]) perturbed by a Wilson loop operator [GorNek2]
@@
    Elliptic Ruijsenaars–Schneider corresponds to hybrid topological-holomorphic $4d$ Chern–Simons ([NekPhD],[Cos13])
In the following we consider the elliptic Calogero?Moser system.
@@ 4 Elliptic Calogero–Moser system
Elliptic Calogero–Moser system ([Cal75]) is a Hamiltonian system containing a triple
@@
One can also consider real-valued $x_{i}$ and $p_{i}$. Analogously to the usual CM this system describes motion of particles interacting with a slightly more sophisticated potential. However allowing positions and momenta to be complex provides an interesting feature: points can swap without gaining infinite energy. This circumstance prompts us to consider an extended phase space.
@@ 4.1 Compactification of the phase space
Instead of giving a full definition, we illustrate the idea in the special case $N=2$. For simplicity, we restrict to the center-of-mass frame
@@ 4.2 Around fixed points
$\begin{cases}X&=\zeta x\\ Y&=\zeta(y+{\mathrm{i}}z)\\ Z&=-\zeta(y-{\mathrm{i}}z)\end{cases}$
@@
$\omega^{\mathbb{C}}=dz_{1}\wedge dz_{2}=\frac{dY\wedge dZ}{4X}={\mathrm{i}}\zeta\frac{dy\wedge dz}{2x}.$
@@
$\langle[\omega^{\mathbb{C}}],[D]\rangle=\langle[\omega^{\mathbb{C}}],[S^{2}]\rangle=\int\limits_{S^{2}}\omega^{\mathbb{C}}={\mathrm{i}}\zeta\int\limits_{S^{2}}\frac{dy\wedge dz}{2x}={\mathrm{i}}\zeta\int\limits_{S^{2}}\mathrm{Vol}_{S^{2}}=4\pi{\mathrm{i}}\zeta.$
@@ 4.3 Phase space as an algebraic variety
5 Complexification of tCMtCM2d2d Yang–Mills correspondence
In the following we describe the complex analogue of the infinite-dimensional Hamiltonian system considered above. The symplectic reduction in this setting will reproduce the elliptic CM system. By complexification we mean passage from left to the right in the following table.
@@ 5 Complexification of tCMtCM2d2d Yang–Mills correspondence
$(g^{-1}Eg,g^{-1}A_{x}g-{\mathrm{i}}g^{-1}\partial_{x}g,g^{-1}(0)z)$ $(g^{-1}Eg,g^{-1}\bar{A}g-g^{-1}\partial_{\bar{z}}g,g^{-1}(0)v,g(0)w)$
@@
$J={\mathrm{i}}(z\otimes z^{\dagger}-\nu 1_{N})$ $J^{\mathbb{C}}={\mathrm{i}}(\nu 1_{N}-v\otimes w^{t})$
@@
([Ati57]) General rank $N$ complex vector bundle of degree 0 on $C_{\tau}$ is isomorphic to $\bigoplus\limits_{i=1}^{N}L_{i}$, where each $L_{i}$ is a degree 0 line bundle.
@@ 5.1 Complex gauging
$g=\mathrm{diag}(e^{\frac{2\pi{\mathrm{i}}}{\tau-\bar{\tau}}((z-\bar{z})n_{i}+(z\bar{\tau}-\bar{z}\tau)m_{i})}):x_{i}\mapsto x_{i}+m_{i}+n_{i}\tau.$
@@ 6 Gauge theories and measures on partitions
Supersymmetric gauge theories in four dimensions, and five and six dimensional theories compactified to four dimensions, produce a class of complexified statistical models. The random variables of these models are collections of Young diagrams, also known as colored partitions [Nek02, Nek15]. More specifically, this is a result of a two-step localization procedure. First of all, it is well-known since the interpretation of M. Atiyah and L. Jeffrey of E. Witten’s topological quantum field theory interpretation [W88b] of Donaldson theory [Don90], path integral of topologically twisted ${\mathcal{N}}=2$ supersymmetric gauge theory, for specific correlation functions, can be interpreted as an integral of equivariant differential forms (albeit in an infinite-dimensional setting). This interpretation has been generalized to theories with matter hypermultiplets [VW94, LM97, MW97, LNS98]. Whenever such interpretation holds, path integral localizes to a finite-dimensional integral over the moduli space of instantons, i.e. solutions to the $F^{+}_{A}=0$ equations modulo suitable gauge group. This localization is an infinite-dimensional version of the equivalence of different representatives of Euler class of a vector bundle: Poincare dual to the zero locus of generic transverse section, and the curvature Gauss-Bonnet type form. The next step undertaken in [Nek02] is to deform the theory in such a way that the path integral measure becomes the equivariant differential form with respect to both the local gauge symmetries and the global symmetries consisting of the rotations of ${\mathbb{R}}^{4}$ (more generally, isometries of the four-manifold on which one studies the theory), the rotations of gauge frame and possibly rotations of the spaces of multiplicities of matter fields. The question of framing arises when one is working on ${\mathbb{R}}^{4}$ or other non-compact manifolds. We need to impose the boundary conditions both on gauge fields and gauge transformations one divides by. As a result, the gauge group $\mathcal{G}$$_{\infty}$ is a proper normal subgroup in the group $\mathcal{G}$} of all gauge transformations defined on some compactification of ${\mathbb{R}}^{4}$, e.g. $S^{4}={\mathbb{R}}^{4}\cup{\infty}$. The quotient ${\mathcal{G}}/{\mathcal{G}}_{\infty}$ is the group of frame rotations. The equivariant parameters $\bf a$ for this group are identified with the vacuum expectation value of the adjoint-valued complex Higgs field, the scalar in the vector multiplet of ${\mathcal{N}}=2$ supersymmetry:
@@
The way path integral localizes onto the instanton configurations and then onto the set of special instanton configurations has been described in literature on many occasions, see [P17] for a review. Using the localization in equivariant cohomology we express the integral as a sum over the corresponding fixed points. The subtlety is that the fixed points do not belong to the actual moduli space of smooth supersymmetric gauge field configurations, such as instantons. Rather, they are parked somewhere on the compactified moduli space. The points of the compactified moduli space no longer parametrize the solutions to the instanton equations on the smooth four dimensional spacetime. They can be interpreted in algebraic geometry as parametrizing torsion free sheaves on ${\mathbb{C}\mathbb{P}}^{2}$ – a compactification of ${\mathbb{C}}^{2}$, a complex plane, whose underlying four dimensional real space is the original spacetime. Every instanton uniquely corresponds to a locally free, hence torsion-free, sheaf. However, there are more general torsion free sheaves. A limit of a sequence of locally free sheaves could be a torsion free sheaf which is not locally free. Let us explain this in a little more detail.
@@
The abovementioned torsion free sheaves obey a weaker condition: if $f\in{\mathcal{O}}_{U}$ is such, that $fs=0$ for non-zero $s\in{\Gamma}({\mathcal{E}}|_{U})$, then $f=0$. However, given the space of local sections of a torsion free sheaf, one cannot, in general, find a $(0,1)$-gauge field $A_{\bar{i}}dz^{\bar{i}}$, so that (6.4) would hold. The issue is that all $\Psi\in{\Gamma}({\mathcal{E}}|_{U})^{\oplus N}$ may degenerate at some point, so that $A_{\bar{i}}$ defined from (6.8) would be singular there. A way to address this issue, by keeping the moduli interpretation of the (partly) compactified moduli space of instantons, is to work on a noncommutative ${\mathbb{R}}^{4}$ [NS98].
@@
in other words, form a partition. Equivariant localization procedure assigns a weight to each fixed point of the form (6.9), and thus defines a measure on a set of multi–partitions. In Sections 6.2–6.3 we introduce some basic notions relevant for partitions and discuss their properties. Then in Sections 6.4–6.8 we give precise formulas for the measures on multi–partitions arising from gauge theories in various examples.
@@
These notions are reviewed in [NS98]. Both (6.13) and (6.14) are invariant under the transformations